OpenStudy (anonymous):

Just a random math question. Is expanding a trinominal a method of factoring?

OpenStudy (anonymous):

yes, if it works

OpenStudy (anonymous):

So in: 9x^2-24x+16, you could factor it by grouping and get (3x-4)(3x-4) or expand it and get (3x^2)-2(3*4)x+4^2?

OpenStudy (anonymous):

@lizzylou169 its the same thing.... if given equation is a perfect square

OpenStudy (anonymous):

Isn't 9x^2-24x+16 a perfect square?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

But what if the function isn't a perfect square? What would I do then?

OpenStudy (anonymous):

okay then depending upon the conditions expand it or factor it...

OpenStudy (anonymous):

Okay then...I'll come up with a random function I suppose. x^2+5x+7

OpenStudy (anonymous):

this function does not have real values..

OpenStudy (anonymous):

Math is frustrating. -__-

OpenStudy (anonymous):

whatever ^-^ tell me you what do you wanna find??

OpenStudy (anonymous):

I just want to see how to factor something by grouping and another method is the function is not a perfect square.

OpenStudy (anonymous):

if*

OpenStudy (anonymous):

let's take a function x^2-5x+4=0 are you comfortable with this one?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

(x^2-5x)-(5x-4) x(x-5) There's no gcf in (5x-4)!

OpenStudy (anonymous):

so now we first expand it by means we modify our equation as x^2-4x-x+4=0 why we did that?? because if you mulitply the last term and first term we get 4x^2 and we want our middle to split in the way so that product of the components of it would be 4x^2 here middle term is -5x (remember sign) so we can write -5x=-4x-x and product of its components(that is of 4x and x ) is 4x^2 this is the method and only applicable to quadratic equation so after expanding in this way we get x^2-4x-x+4=0 now take x as a common then we get x(x-4)-x-4=0 then take x-4 as common we get (x-1)(x-4)=0 so x-1=0 or x-4=0 so x=1 or x=4

OpenStudy (anonymous):

@lizzylou169 hope it helps..

OpenStudy (anonymous):

Thank you. :)

OpenStudy (anonymous):

you are most welcome...!!